Your groceries are in a bag with paper handles. The handles will tear off if a force greater than 53 N is applied to thenm (a) What is the greatest mass of groceries that can be lifted safely with this bag, given that the bag is raised with constant speed? kg (b) What if it is raised with an acceleration of 1.45 m/s ? kg
Answer #1

Fap app m 53 = m(9.8) Q 145 ap- 53 = nn (1.45+3.8 53 =g-72k 1.25